OpenStudy (anonymous):

Solve using the quadratic formula: -r^2-5r+1=0

hero (hero):

Just use the quadratic formula

OpenStudy (anonymous):

I get to \[r=5\pm \sqrt{21}/-2\] Am I on the right track?

hero (hero):

No

OpenStudy (anonymous):

Ugh

hero (hero):

You must have made a mistake somewhere

hero (hero):

I think you subtracted when you should have added

hero (hero):

Oh, I know what you did wrong

hero (hero):

re-write it like this first: r^2+5r - 1 = 0

OpenStudy (anonymous):

So, \[r=5\pm \sqrt{(-5)}^{2}-4 * 1 * 1 / -2\]

OpenStudy (pokemon23):

Hero wish me luck going to a chess tournament

OpenStudy (anonymous):

You'll get \[r =5\pm \sqrt{29}\div -2\]

hero (hero):

LUCK

OpenStudy (anonymous):

\[Ax^2 + Bx + C = 0\] \[x = \frac{-B \pm \sqrt{B^2-4\cdot A \cdot C}}{2\cdot A} \] is the quadratic formula...

OpenStudy (anonymous):

yeah, so i ended up with 25 - 4 / -2 Should I have added?

OpenStudy (anonymous):

\[-r^2 - 5r + 1 = 0\] A = -1, B = -5, C = 1, so \[x = \frac{5 \pm \sqrt{25+4}}{-2} = \frac{5 \pm \sqrt{29}}{-2}\]

OpenStudy (anonymous):

jermurray how do you make a proper divided by line like so?

OpenStudy (anonymous):

This system understands LaTeX. You can find the syntax in any basic LaTeX tutorial. In particular, you start an expression with \[, you end it with

OpenStudy (anonymous):

\], and in the middle goes \frac{3}{4} if you want to type \[ \frac{3}{4} \]